Math

Find the center, radius and write the standard form: x2 + y2 + 6x ­ 10y + 10 = 0                    
Answer #1

x2 + 6x + y2 ­ 10y = ­10

x2 + 6x + 9 + y2 ­ 10y + 25 = ­10 + 9 + 25

 (x + 3)2 + (y ­ 5)2 = 24

Center: (­3, 5) r = √24


answered by: Tina Gao